40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 193 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  7386   Fri Sep 14 01:35:55 2012 Mike J.UpdateElectronicsAS beam scan PLOTS

H_razor.jpegV_razor.jpeg

  7385   Fri Sep 14 01:18:51 2012 ranaUpdateCOC2 Layout Changes

After looking at the in-vacuum layout I think we should make two changes during the next vent:

1) Reduce the number of mirrors between the FI and its camera. We install a large silvered mirror in the vacuum flange which holds the Faraday cam (in the inside of the viewport). That points directly at the input to the Faraday. We get to remove all of the steering mirror junk on the IO stack.

2) Take the Faraday output (IFO REFL) out onto the little table holding the BS and PRM Oplevs. We then relocate all 4 of the REFL RFPDs as well as the REFL OSA and the REFL camera onto this table. This will reduce the path length from the FI REFL port to the diodes and reduce the beam clutter on the AS table.

  7384   Fri Sep 14 01:05:36 2012 DenUpdateSUSMirrors being installed on active TTs

Quote:

I have given Den 4 G&H R>99.99% mirrors to be installed on the 4 active tip tilts. 

 I've installed the mirrors on 4 tip-tilts. I was able to align 3 of them in pitch, the last one has a screw with damaged thread, I'll continue with it tomorrow.

Alignment accuracy in pitch is ~0.1 mrad. Mirrors oscillate a lot probably due to air flow coming from the side wall.

  7383   Fri Sep 14 00:56:13 2012 JenneUpdateGeneralDRMI aligned

[Rana, Jenne]

We aligned the DRMI, and have concluded that it looks good enough that we should close up and pump down soon. We still need to use the camera to check things, and get all pickoff beams out of the chambers, so don't get too excited yet.

We looked at the mode matching telescope's calculated beam propagation, and since we're using spherical telescope mirrors at non-zero degree incidence angle, we expect an astigmatism about like what we are seeing on the AS camera.  This matches up with the measurements that Mike posted from his and Q's measurements earlier today.  We think that it has 'always' been this way, and someone just picked a camera position such that the beam used to look more round than it does now.

We aren't entirely sure what's up with the SRM - it almost looks like the pitch and yaw are coupled, but it was pretty easy to align the PRMI.  We don't see any evidence of the crazy, crappy beam that we did before the vent.  This means we have fixed most of the bad clipping problems we were seeing over the last ~year.

In the process of aligning the DRMI, we fixed up the input beam alignment - we were not hitting the exact centers of the MMT mirrors (in pitch, mostly), so we fixed that, and propagated the alignment fix through the chain.  In all, we touched the knobs on PZT1, MMT1, MMT2, PZT2.  The beam then went through the SRM, and we touched a few of the output steering mirrors to get the beam centered on all mirrors. 

I remeasured the MC spot positions, and they're a little worse than they have been.  Some of the spots seem to be off by 1.75mm (or less) on MC 1 and 3.  The numbers, MC1,2,3 pitch, then MC1,2,3 yaw are:   1.749759        9.744013        1.025681        -0.791683       -1.338786       -1.779958

A question to consider before doing the final-final alignment checking is: do we need to get the MC spots centered better than this, especially in light of the potential PMC axis having moved? 

  7382   Fri Sep 14 00:33:31 2012 ranaUpdatePSLPMC alignment - mystery in reflected power

The PMC reflection made it seem that the beam going into it was misaligned. I went to the table and aligned the input beam to maximize the PMC transmission. I got ~10% improvement.

Just to check if something was loose, I started tapping things upstream of the Faraday. When I tapped the actual PMC body it seemed to get unseated and the reflected (unlocked) power jumped up by more than a factor of two.

I don't understand how this could be. The attached trend of the PMC channels shows that ever since the PSL upgrade, the PMC refl has been at the low level of ~0.3 V, except for a brief phase soon after the upgrade late in 2010 and then also for a few hours early in May of 2012.

If the PMC body actually moved, it seems that the pointing into the MC would also change and I don't see that. So what else can it be? Is there some clipping or dust or a burn spot on the PMC REFL path?

The PMC refl image was lost after the body re-settled itself. Jenne and I re-aligned it and added a 0.5 ND filter to the existing ND in order to account for the higher power. We should hide all of the reflective ND filters and just use absorbtive ND for the cameras to prevent reflections.

a.png

This image of the past hour shows the event at just before midnight (0650 UTC) where the PMC reflection goes up from 0.28 to 0.85.

Attachment 1: pmcr.jpg
pmcr.jpg
Attachment 2: e.png
e.png
  7381   Thu Sep 13 23:27:14 2012 JenneUpdateGeneralPre-close checklist

We need to do the following things:  Images of optics in DRMI chain, place black glass beam dumps, make sure pickoff beams get out, align IP POS/ANG.

Black glass: behind MMT1, behind IPPOSSM3, forward-going POP beam.

Images and pickoff stuff should happen at the end of each vent.

Images need to be taken of the following optics (with ruler edge at center of optic):

* PZT1

* MMT1

* MMT2

* PZT2

* PRM

* PR2

* PR3

* BS (front and back?)

* ITMX

* ITMY

* SR3

* SR2

* SRM

* OM1

* OM2

* OM3

* OM4=PZT3

* OM5=PZT4

* OMPO

* OM6

* Viewport as AS beam leaves chamber

* POYM1 (check no clipping on edge of mount)

* POXM1 (check no clipping on edge of mount)

Pickoff / aux beams:

* REFL path

* POX

* POY

* POP

* IPPOS

* IPANG

  7380   Thu Sep 13 19:59:43 2012 Mike J.UpdateElectronicsAS beam scan

**EDIT:** Mixed up X and Y. Beam is 3.5844 mm tall and 2.7642 mm wide

14.112 hundredths of an inch in the vertical direction

3.5844 millimeters

10.883 hundredths of an inch in the horizontal direction

2.7642 millimeters

Plots and error bars to come soon.

  7379   Thu Sep 13 17:19:45 2012 JenneUpdateSUSMirrors being installed on active TTs

I have given Den 4 G&H R>99.99% mirrors to be installed on the 4 active tip tilts.  He's in there working on things (incl. installing and balancing the pitch of the mirrors) right now.  He'll elog his work later.

  7378   Thu Sep 13 07:36:41 2012 SteveUpdateGeneralrestore conditions

Quote:

 Summary: Recorded the presence of higher order modes in IMC

What I did: Misaligned the flat mirror MC1 by small amount in both pitch and yaw (it was needed to be done cause at the beginning of the experiment no higher order modes were present)  and scanned the cavity for frequency-range 32MHz to 45MHz.

I found the presence of higher order modes around 36.7MHz (1st order)  and 40.6MHz (2nd order) along with two other strong modes near 35MHz and 42.5MHz.

 

 Please, restore condition after you finished and update elog right away! People wasted hours yesterday not knowing the condition of the MC

  7377   Wed Sep 12 20:08:51 2012 ericqUpdateElectronicsAS beam scan

Quote:

We conducted a beam scan on the AP table of the AS beam. We used a lens to focus the beam onto a power meter, and slowly moved a razor blade across the beam using a micrometer, vertically and horizontally both in front of and behind the beam. We also had to block the beam next to the AS beam in order to do this, but is unblocked now. Mike will begin curve fitting the data to try and see if there is a different spot size given by the x-axis vs. the y-axis, and if the lens has any effect.

 [ericq, mikej, some input from zach]

After realigning the MC, the measurement was repeated this afternoon. This time, however, we isolated the beam from ITMY by misaligning ITMX. The beam looked somewhat elliptical to me, and Mike should have fits up tonight. Afterwards, ITMX was returned to the position I found it in, and the PMC shutter and access connector were closed. (Sorry about last night!)

  7376   Wed Sep 12 19:26:08 2012 Rijuparna ChakrabortyUpdate cavitymode scan

 Summary: Recorded the presence of higher order modes in IMC

What I did: Misaligned the flat mirror MC1 by small amount in both pitch and yaw (it was needed to be done cause at the beginning of the experiment no higher order modes were present)  and scanned the cavity for frequency-range 32MHz to 45MHz.

I found the presence of higher order modes around 36.7MHz (1st order)  and 40.6MHz (2nd order) along with two other strong modes near 35MHz and 42.5MHz.

 

Attachment 1: P120912_11.32.jpg
P120912_11.32.jpg
Attachment 2: P120912_14.13.jpg
P120912_14.13.jpg
Attachment 3: P120912_14.17.jpg
P120912_14.17.jpg
Attachment 4: P120912_11.25.jpg
P120912_11.25.jpg
Attachment 5: P120912_14.09.jpg
P120912_14.09.jpg
Attachment 6: P120912_14.30.jpg
P120912_14.30.jpg
Attachment 7: P120912_14.34.jpg
P120912_14.34.jpg
  7375   Wed Sep 12 17:02:00 2012 SteveUpdateCamerasbaffle plate hole getting larger

Quote:

Quote:

The alignment of the pick-off mirror near ETMX is done. Everything turned out to be easy once we realized that there is no sense getting the alignment laser (going through viewport to pick-off to ITMX) back to ETMX. It is only necessary to hit ITMX somehow, since this makes sure that there is one scattered beam that will make it from ITMX to pick-off through viewport.

After the auxiliary optic (that we never used in the end) was removed again, we levelled the optical table.

So in the current setup, we can have small-angle scattering measurements on ITMX and large-angle scattering measurements on ETMX.

 This is how it was envisioned. The video camera was in nobodies mind to look through the 40 mm  diameter hole than.

 Rana is proposing 50 mm hole in the baffle plate that is attached to the tower.  Atm1

Atm2 is showing the back side where the solid line is 40 mm

Attachment 1: IMG_1631.JPG
IMG_1631.JPG
Attachment 2: IMG_1628.JPG
IMG_1628.JPG
  7374   Wed Sep 12 11:33:49 2012 SteveUpdateSAFETYsafety training

Eric Quintero and Mike Jenson received 40m specific basic safety training.

  7373   Wed Sep 12 08:16:49 2012 SteveConfigurationPEMchamber must be sealed overnight!

Quote:

We conducted a beam scan on the AP table of the AS beam. We used a lens to focus the beam onto a power meter, and slowly moved a razor blade across the beam using a micrometer, vertically and horizontally both in front of and behind the beam. We also had to block the beam next to the AS beam in order to do this, but is unblocked now. Mike will begin curve fitting the data to try and see if there is a different spot size given by the x-axis vs. the y-axis, and if the lens has any effect.

 The vacuum envelope must be sealed with light doors on o-rings to insure a bug free IFO.  This was a violation!

  7372   Tue Sep 11 17:17:51 2012 Eric Q., Mike J.ConfigurationElectronicsAS beam scan

We conducted a beam scan on the AP table of the AS beam. We used a lens to focus the beam onto a power meter, and slowly moved a razor blade across the beam using a micrometer, vertically and horizontally both in front of and behind the beam. We also had to block the beam next to the AS beam in order to do this, but is unblocked now. Mike will begin curve fitting the data to try and see if there is a different spot size given by the x-axis vs. the y-axis, and if the lens has any effect.

  7371   Mon Sep 10 19:04:32 2012 ranaUpdateGeneralplan

 On Friday, Koji and I adjusted the beam pointing into the DRMI using the PZT yaw and found that the beam inside the DRMI (as seen on the AS camera) looked OK (not distorted too much).

So it seems that the issue seen before, namely that the DRMI resonant mode is very strange, is no longer true.

The camera image at the AS port still looks elliptical. So Jenne and Mike have started to make this beam round by adjusting the lenses.

Our plan now is:

1) Fix AS camera optics to get a round beam (single bounce off of ITMY).

2) Flash DRMI to make sure the beam at AS is still round.

3) Using the moveable Watec camera and Sensoray, get images of the spot on all DRMI mirrors with DRMI flashing. Use targets and rulers whenever possible to get quantitative measurements of the beam positions. (i.e. just saying "Oh, its pretty much in the center" is the Mickey Mouse approach to science)

4) Align all pickoff beams in this situation. Make sure there is no in vac clipping. Align IP POS and ANG using this input beam pointing.

5) Pump down.

 

  7370   Mon Sep 10 18:42:33 2012 Jenne, Mike J.UpdateCamerasXY beam scan tomorrow

We tweaked the mirror on the AP table to go through the center of the lens in order to get a more circular beam, but it seemed ineffective. So we put an IR card in front of the lens and behind the lens to see if the beam was circular or ovacular, but could not tell. We also moved the camera to see, but still couldn't see a distinct circle or oval. So Mike and Q will do a beam scan tomorrow in both the X and Y directions to see if the beam is circular or not.

  7369   Mon Sep 10 08:50:35 2012 SteveUpdateGeneralBeam scan measurement plan - to do Monday morning.

 

 I misaligned ITMX pitch on Friday and brought out the beam at 44" height. The beam was bouncing to much. I only realized it this morning why. The OSEM voltages are 1.8, 1.7, 0.2 and 0.9V  Even with a stable 8-9 mm diameter beam you would be clipping

on the beam scanner 9 mm aperture. You can bring out the beam with one mirror right after  PR3, just remove  PRMOP2

  7368   Sat Sep 8 00:15:57 2012 Rijuparna ChakrabortyUpdate cavitymode scan

Quote:

Quote:

 IMC transmission photodiode has been aligned.

 Which PD?  The 'regular' DC one, or the newer one?  Why did it need realigning?  What mirrors did you touch to do the alignment?

Did you do anything else in the last 3 days?  I want to see ALL the gory details, because it can help people doing future measurements, or help us debug if something is wrong with the interferometer later.

MORE WORDS! Thanks.

 No, not the "regular DC one", the "newer one"  along with the controls of the corresponding mirror only i touched.

It needed to be realigned cause last week when we fitted a longer cable there, which may reach the network analyzer, it got misaligned since it got touched.

No other component in that box except that PD and the corresponding mirror controls I touched.

For my last 2 days work, I feel my last elog is reliable.

Today other than doing this, I checked for the higher order modes of the cavity, misaligning one of the MC mirror though the software only. I didn't mention it in my elog cause although I saw the presence of the higher order modes I didn't record it, so I can not upload any picture in support of such a statement.

Thanks

  7367   Sat Sep 8 00:04:53 2012 JenneUpdateGeneralBeam scan measurement plan - to do Monday morning.

[MikeJ, Jenne]

We have a plan for how we're going to measure the beam after PR3.  Mike is going to write up a nifty program that will spit out the waist of the beam if you give it a bunch of razor blade measurement data.

Since the beam bounced off of the pitched ITMX is coming out of the chamber so high, it's kind of a pain to setup optics to steer the beam down the walkway next to the Yarm.  So, I have a new vision.

I think that we can get the beam right after PR3 onto the PRM/BS oplev table using 3 clean mirrors (of which we have many spares, already clean).  Once on the oplev table, we can put a 2" Y1 mirror to steer the beam down the walkway, after taking off the short east side of the table.  Then we can use the little breadboard on the mobile blue pedestal for the razor blade / power meter setup.

The razor blade on a micrometer translation stage will be the first thing on that table that the beam sees. Then, a 2" lens to get the beam small enough to fit on the power meter.  Then, obviously, the power meter.  We can measure the distance between the oplev table and the razor blade using the laser range finder, which has pretty good accuracy (it's sub-centimeter, but I don't remember the exact number for the precision).

A lens is not okay if we're trying to get the beam directly onto the beam scanner, since it will distort the beam.  However, as long as the razor blade is before the lens, and we're just using the lens to get the full intensity of the non-obscured part of the beam onto the power meter, I think using a lens should be fine.  If we don't / can't use a lens, we're going to run into the same problem we have with the beam scanner, since the power meters all have a fairly small aperture.  Even the big 30W power meter's aperture will be on the order of the size of the beam, so we won't be able to guarantee non-clippage.

The main problem I see with the technique as I have described it, is that the beam is going to hit 4 mirrors (3 in-vac, one outside) before going to the razor/lens/power meter.  We have to make sure that we're not clipping on any of those mirrors.  Also, this measurement version takes the beam after PRM, PR2 and PR3, but not after the BS and ITM.  I don't think we're concerned with either of those 2 optics, (especially since this is refl off the front of the BS, so won't see any potential clipping on the BS cage), but just in case we are, this measurement isn't so useful, and we'd have to come up with a different way of placing the mirrors on the in-vac tables to get a beam bounced off  of  a yaw-ed ITMX. 

Perhaps it would be easier to just go with the pitched ITMX version of the measurement, but I could use some ideas / advice on how to mount mirrors and lenses ~4 feet off the ground outside of the chambers, and not have them waving around on skinny sticks.

 

EDIT: Another idea is to instead use the beam transmitted through the BS, put a single clean steering mirror in the ITMY chamber, and get the beam out of the ITMY door.  This could either be the beam before the ITM, or we could yaw the ITM a little and take the reflected beam.

  7366   Fri Sep 7 17:37:16 2012 JenneUpdate cavitymode scan

Quote:

 IMC transmission photodiode has been aligned.

 Which PD?  The 'regular' DC one, or the newer one?  Why did it need realigning?  What mirrors did you touch to do the alignment?

Did you do anything else in the last 3 days?  I want to see ALL the gory details, because it can help people doing future measurements, or help us debug if something is wrong with the interferometer later.

MORE WORDS! Thanks.

  7365   Fri Sep 7 17:34:53 2012 JenneUpdateComputersSensoray Video Capture

Quote:

To capture video with the Sensoray, open the GUI (python ./demo.py), simply press "Save," enter a filename, and hit "Stop" when you wish to stop recording. If you want to change the video format, there is a dropdown menu labelled "Format." I recommend MP4 for standard video, and nv12 for RAW video.

 I also installed mplayer on rossa, so we can play the videos there.

Even though Mike won't admit it, the video stuff is all in /users/sensoray/ .  I opened the demo.py from there, and it also works.

  7364   Fri Sep 7 17:24:16 2012 Mike J.UpdateComputersSensoray Video Capture

To capture video with the Sensoray, open the GUI (python ./demo.py), simply press "Save," enter a filename, and hit "Stop" when you wish to stop recording. If you want to change the video format, there is a dropdown menu labelled "Format." I recommend MP4 for standard video, and nv12 for RAW video.

  7363   Fri Sep 7 15:58:29 2012 Rijuparna ChakrabortyUpdate cavitymode scan

 IMC transmission photodiode has been aligned.

  7362   Fri Sep 7 15:31:52 2012 Mike J.UpdateComputersSensoray back up

Video Capture with the Sensoray works again. Pianosa just needed mplayer installed for it to play properly.

Attachment 1: output_5.mp4
  7361   Fri Sep 7 13:01:53 2012 ManasaConfiguration40m UpgradingBaffle problem

Quote:

Quote:

The required diameter for the baffle if it sits on the cage at 1.77" from the test masses: the current baffle (dia. 40mm) centered along the beamline, will allow ~8.6mm visibility from the center of the test mass (in case of ETMY).

*assuming the pick off mirror is placed at the edge of the tunnel

Estimations of the visibility region (r1 on the test mass) with baffle (aperture size 40mm).

The baffle is installed on the cage at 1.125" from the test mass (distance changed from the previous elog after a double check).

The 40mm aperture is in no way going to help get clear view of the ITMs; 

Required baffle diameter to have a visibility region r1 = 3 times the beam diameter

Picture1.png

  7360   Fri Sep 7 12:28:09 2012 KojiUpdateLSC11&55MHz modulations turned off

11MHz modulation source was turned off (disabled) at Marconi at 12:00.

  7359   Fri Sep 7 11:58:12 2012 ManasaConfiguration40m UpgradingBaffle problem

Quote:

The required diameter for the baffle if it sits on the cage at 1.77" from the test masses: the current baffle (dia. 40mm) centered along the beamline, will allow ~8.6mm visibility from the center of the test mass (in case of ETMY).

*assuming the pick off mirror is placed at the edge of the tunnel

Estimations of the visibility region (r1 on the test mass) with baffle (aperture size 40mm).

The baffle is installed on the cage at 1.125" from the test mass (distance changed from the previous elog after a double check).

The 40mm aperture is in no way going to help get clear view of the ITMs; 

Attachment 1: bfl.png
bfl.png
  7358   Fri Sep 7 09:37:20 2012 SteveUpdateCamerasbaffle plate for SOS

Quote:

The alignment of the pick-off mirror near ETMX is done. Everything turned out to be easy once we realized that there is no sense getting the alignment laser (going through viewport to pick-off to ITMX) back to ETMX. It is only necessary to hit ITMX somehow, since this makes sure that there is one scattered beam that will make it from ITMX to pick-off through viewport.

After the auxiliary optic (that we never used in the end) was removed again, we levelled the optical table.

So in the current setup, we can have small-angle scattering measurements on ITMX and large-angle scattering measurements on ETMX.

 This is how it was envisioned. The video camera was in nobodies mind to look through the 40 mm  diameter hole than.

Attachment 1: IMG_1624.JPG
IMG_1624.JPG
Attachment 2: IMG_1618.JPG
IMG_1618.JPG
Attachment 3: IMG_1616.JPG
IMG_1616.JPG
  7357   Fri Sep 7 01:25:53 2012 JenneUpdateGeneralPRC, SRC flashing

[Koji, Jenne]

* Found that IPANG was no longer centered, so we used PZT2's sliders to get the spot back on the center of the QPD.  Koji points out that I should have moved the lens even farther away, to have a larger beam (many mm, not just ~1) on the QPD.

* Found that MICH alignment had drifted, so used ITMX to realign MICH.

* Aligned PRM, got REFL beam through viewport.  Just made sure reflected beam was colinear with incident beam.

* PRC flashes were visible on AS camera. 

* PRM was more precisely aligned to have good interference with ITM reflections, by looking at AS camera.

* Decided to align SRM.  Spot was ~5mm too far to the north on the SRM....so we were off from center by ~5mm.

* Moved SR2 yaw a little bit to get spot centered on SRM.

*  Couldn't align SRM within bias slider range, so moved SRM in yaw to get reflected beam colinear with incident beam.

* Centered the spot on the steering mirrors.  The 2nd steering mirror after the SRM was moved by ~1 inch.  All mirrors after that were aligned to match this new beam.

* Found spot on AS table, aligned AS table mirrors so that beam hits AS55 PD window.  Haven't actually centered beam on PD.

* Transmission of 99% reflector was too weak to use with a card to get the beam back on the AS camera, so we moved the camera over to the AS110 path.

* Precisely aligned PRM and SRM by watching AS camera.

* Both the PRC and SRC look kind of funny.  Koji agrees.  Seriously.  They're a little weird. We can't align either recycling cavity, one ITM at a time (so PRM with ITMX, PRM with ITMY, SRM with either single ITM) to get rid of all the fringes.  Something is definitely funny.  It's got to be in the recycling cavities, since the weirdness is common between both ITMs for a given recycling mirror.  We need to take Sensoray views of these tomorrow.=

* There is some clipping on the right side of the AS camera view.  We have determined that it is not clipping at the viewport exiting the vacuum, but we aren't sure where it is.  It is at least before PZT4 (the 2nd PZT in the output AS path). 

  7356   Fri Sep 7 00:08:10 2012 janoschMetaphysics baffle clipping loss

With a curvature radius of about 57m for the ETMs, flat ITMs at the beam waist, and using 39m for the arm lengths, one finds that the beam radius at the ETMs is about 5.3mm. The clipping power loss of a 5.3mm beam through a 20mm radius baffle hole would be less than a ppm of a ppm if the beam was perfectly centered. If the baffle hole had 15mm radius, the clipping loss would be 0.01ppm. If the baffle hole had 10mm radius, the loss would be 810ppm. The loss values are calculated using the formula of the  "Gaussian beam" Wikipedia article, "Power through an aperture" section. So I did not check if that one is ok.

  7355   Thu Sep 6 19:36:19 2012 JenneUpdateRF SystemAS 55 is fine

Quote:

I was going to lock MICH, but I don't see anything on dataviewer for either AS55Q or ASDC.  I went out onto the table, and there is beam on the diode, but no mV out on a voltmeter connected to the DC monitor point.  I shine a flashlight, and still I see 0.0mV.  So, something is up with AS55, but since the michelson is aligned right now, I'm not going to mess with the PD.  I won't lock MICH, I'll just move on.  Koji is taking a look at the diode, but if he doesn't get it figured out tonight, we can take a closer look after we pump down.

 Never mind.  I was using an LED flashlight, which doesn't emit light that the PD is sensitive to.  A regular flashlight gives plenty of signal on the DC out. 

Using an SR560 with 30Hz low pass and gain of 100, it was pretty easy to align the light on the PD. 

Koji calculates in his head that there is about 6 microwatts of light incident on the PD, which is not a lot of light. Our SNR may be kind of lame for locking right now.

  7354   Thu Sep 6 19:21:58 2012 ManasaConfiguration40m UpgradingBaffle problem

For the current baffle (dia. 40mm) centered along the beamline place at 1.77" from the test mass, the baffle will allow ~8.6mm visibility on the camera from the center of the test mass (in case of ETMY).

*assuming the pick off mirror is placed at the edge of the tunnel

Attachment 1: bfl.png
bfl.png
  7353   Thu Sep 6 18:49:30 2012 JenneUpdateRF SystemAS 55 may be broken

I was going to lock MICH, but I don't see anything on dataviewer for either AS55Q or ASDC.  I went out onto the table, and there is beam on the diode, but no mV out on a voltmeter connected to the DC monitor point.  I shine a flashlight, and still I see 0.0mV.  So, something is up with AS55, but since the michelson is aligned right now, I'm not going to mess with the PD.  I won't lock MICH, I'll just move on.  Koji is taking a look at the diode, but if he doesn't get it figured out tonight, we can take a closer look after we pump down.

  7352   Thu Sep 6 17:11:40 2012 janosch, Manasa, SteveUpdate pick-off and baffle at ETMY

We have installed the pick-off mirror at the ETMY table for the small-angle scattering measurement on ITMY. As we had already done for the X arm pick-off, the pick-off mirror at ETMY was aligned shooting a green laser normally through the viewport on the pick-off and steering it onto ITMY.

A baffle was also installed at a distance of about 30cm from ETMY near the edge of the table.

  7351   Thu Sep 6 17:06:25 2012 Rijuparna ChakrabortyConfigurationelogCavitymode scan

 Aim: to scan the cavitymodes of IMC

The circuit used: 

 Attachment4

Results obtained:

Attachment 1,2,3

 

Attachment 1: 3.pdf
3.pdf
Attachment 2: 2.pdf
2.pdf
Attachment 3: 1.pdf
1.pdf
Attachment 4: cavityscanconnections.pdf
cavityscanconnections.pdf
  7350   Thu Sep 6 16:46:44 2012 JenneUpdateSUSBS aligned, target removed

Q and I aligned the BS such that we were hitting the center of ETMX. The ETMX cage does not have OSEM setscrew holes on the front, so it is not possible to put the targets that Steve made on this optic.  So, I put the freestanding ruler in front of the optic, with the edge of the ruler at the center (as viewed from above) of the optic.  Then Eric steered the BS until we were hitting the 5.5" mark, and roughly half of the beam was obscured by the ruler.

We then aligned ITMX such that the prompt reflection was colinear with the incoming beam. 

I checked the 2 spots through the BS, heading to the AS port.  (2 spots since MICH hasn't been locked / finely aligned yet).  They were being clipped on the 2nd output PZT.  I adjusted the knobs of the first output PZT to center the spots on the 2nd PZT.  Note that the output PZTs' power is still off, and has been off for some unknown length of time.  I had found them off when prepping for the vent a week or two ago.  So the current alignment depends on them staying off.  We don't really need them on until we're ready to employ our OMC.

The beams now look nicely unclipped on the AS camera, and we're aligning MICH.

  7349   Thu Sep 6 13:07:02 2012 JenneUpdateIOOIPANG no longer a reference :(

I was having trouble centering IPANG using the PZTs, and I suspected something funny was going on at the end.  I went down there, and the beam was focused right on the PD, and the spot was very very small.  I think this means that when I was trying to center the beam, I was falling into the gap between the pieces of the diode.  Also, as Koji pointed out to me the other day, if the PD is at the focal point of the beam, any parallel rays hitting the lens just before the PD will all go to the same place, no matter how the input beam has moved.  This means we're not getting as much info out as we'd like.

So.  I moved the lens a little bit farther from the PD such that we are just beyond the focal point of the beam.  The beam size is now ~1mm on the QPD.

This means, however, that I moved the beam on the QPD such that IPANG is no longer a reference of the input pointing. Ooops. I think this adjustment needed to be done though.  Right now, the PZTs are set to where we had them yesterday, when we moved them slightly to center the IPANG QPD, and I've recentered IPANG.

  7348   Thu Sep 6 10:57:27 2012 JenneUpdateGeneralForgot to turn green refl pd back on

Quote:

I couldn't understand the Y-End green setup as the PD was turned off and the sign of the servo was flipped. Once they are fixed, I could lock the cavity with the green beams.

Quote:

[EricQ, Jenne, brains of other people]

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

 

 I had turned the green refl PD off on Tuesday while we were doing the IPANG alignment, since the beam was not so bright, and the LED on top of the PD was very annoyingly bright.  I forgot to turn it back on.  The sign flip on the servo, I can't explain.

  7347   Thu Sep 6 02:51:47 2012 KojiUpdateGeneralGreen beam roughtly aligned (Re: Yarm aligned to IR incident beam)

The Y-End green beam was roughly aligned by the steering mirrors for the green beam.

I couldn't understand the Y-End green setup as the PD was turned off and the sign of the servo was flipped. Once they are fixed, I could lock the cavity with the green beams.

After a long alignment session, TEM00 was found. The alignment of the green beam has not been optimized.

Looking at the spot position at ETMY OSEM holders (not by the ccd image), it seems that the cavity mode is not at the center of the mirrors.

Quote:

[EricQ, Jenne, brains of other people]

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

 

Attachment 1: green_y.png
green_y.png
  7346   Wed Sep 5 19:29:45 2012 JenneUpdateGeneralYarm aligned to IR incident beam

[EricQ, Jenne, brains of other people]

Checked at ETMY that the pointing hadn't changed a whole lot.  Jamie and Koji pointed out that we were half falling off of the IPANG QPD.  Adjusted PZT2 sliders so that we were again centered on IPANG's QPD.  Before we close up, we'll want to put the sliders back to their nominal positions, and use the knobs to hit IPANG, but this is equivalent and fine for now.  The tip tilts don't seem to have moved much overnight, since the beam drift on both IPANG and ETMY was fixed simultaneously with PZT2 (recall, IPANG picked off before tip tilts exist in the beam path).  This left us hitting the center of ETMY.  We moved ETMY sliders to make the reflected beam hit the center of ITMY (same spot position as transmitted beam from BS).  Then moved ITMY to get prompt reflection to hit same spot on ETMY as original primary beam from BS.  Checked at ITMY that we didn't need to move ETMY anymore.  (Actually, I forget how many iterations we did, but in the end, all of the reflections that we can find are co-located on the test masses.)

Next up:

Align BS so we're hitting the center of ETMX

Tap / readjust ITMX OSEM which is at 0.3 to get it back to the center of its range

Align ITMX to lock MICH

Check no clipping on POX / POY optics, no clipping around BS

Check PRM, SRM alignment (what exactly do we want to do here? Try to lock PRMI / SRMI?)

Get IPPOS out of vac

Fix clipped ITMY / SRM oplev

Install 'black' glass beam dumps - forward-going POP beam, 2 places in BS chamber (check old elog from Jenne/Yuta for the places).

Get green spots co-located with IR spots on ETMs, ITMs, check path of leakage through the arms, make sure both greens get out to PSL table

  7345   Wed Sep 5 13:11:43 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

Quote:

Quote:

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

 Can we see the full design?  If we can't lock the mode cleaner with this thing on then it's really of no use.  We want it to be equivalent to the light doors, but allow us to keep the mode cleaner locked.  That's the most important aspect.

 It also needs to be wide enough that the MMT beam can go through, so that we can not only lock the MC, but also work on the rest of the IFO.

  7344   Wed Sep 5 10:50:15 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

Quote:

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

 Can we see the full design?  If we can't lock the mode cleaner with this thing on then it's really of no use.  We want it to be equivalent to the light doors, but allow us to keep the mode cleaner locked.  That's the most important aspect.

  7343   Wed Sep 5 09:50:25 2012 SteveOmnistructureVACbetter in-air "lite" access connector needed

Quote:

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

 It is in the shop. It will be ready for the next vent. Koji's dream comes through.

Attachment 1: IMG_1612.JPG
IMG_1612.JPG
  7342   Tue Sep 4 20:25:22 2012 jamieOmnistructureVACbetter in-air "lite" access connector needed

We really need something better to replace the access connector when we're at air.  This tin foil tunnel crap is dumb.  We can't do any locking in the evening after we've put on the light doors.  We need something that we can put in place of the access connector that allows us access to the OMC and IOO tables, while still allowing IMC locking, and can be left in place at night.

  7341   Tue Sep 4 20:20:47 2012 jamieUpdateGeneralproblematic tip-tilts

Quote:

We clearly need a better plan for adjusting the tip tilts in pitch, because utilizing their hysteresis is ridiculous.  Koji and Steve are thinking up a set of options, but so far it seems as though all of those options should wait for our next "big" vent.  So for now, we have just done alignment by poking the tip tilt.

Tomorrow, we want to open up the MC doors, open up ETMY, and look to see where the beam is on the optic.  I am concerned that the hysteresis will relax over a long ( >1hour ) time scale, and we'll loose our pointing.  After that, we should touch the table enough to trip the BS, PRM optics, since Koji is concerned that perhaps the tip tilt will move in an earthquake.  Jamie mentioned that he had to poke the tip tilt a pretty reasonable amount to get it to change a noticeable amount at ETMY, so we suspect that an earthquake won't be a problem, but we will check anyway.

 I'm very unhappy with the tip-tilts right now.  The amount of hysteresis is ridiculous.  I have no confidence that they will stay pointing wherever we point them.  It's true I poked the top more than it would normally move, but I don't actually believe it wouldn't move in an earthquake.  Given how much hysteresis we're seeing, I expect it will just drift on it's own and we'll loose good pointing again.

And as a reminder, IPPOS/ANG don't help us here before the tip-tilts are in the PRC after the IP pointing sensors.

I think we need to look seriously at possible solutions to eliminate or at least reduce the hysteresis, by either adding weight, or thinner wire, or something.

  7340   Tue Sep 4 20:13:46 2012 JenneUpdateGeneralThe Plan - Tues evening version

 Tues

* Hit center of ETMY, using input optics, PR3.

* Get IPANG out of vac, center QPD.

Wed

* AM: Riju do MC mode scans

* Starting right after 40m meeting, if not before: Lock MICH to get BS, ITMs aligned well

* Check if beam is hitting center of ITMX.

* Check for clipping around BS

     - Use Watek to look at beam at all 4 BS ports - make sure no clipping going into BS, after BS in the michelson, or the AS port

       - Use some old in-vac mirrors to direct beam out the BS door.  Cameras are waiting near BS chamber.

* Prepare glass beam dumps??

* IPPOS - make sure beam gets out of chamber

* Jan take photos of ETMX scattering setup

* Manasa take in-vac photos of all tables, for table layouts

* Jan / Manasa viewport transmission

* Install glass beam dumps?

* Install glass baffle at ETMY. Jan maybe install baffle at one of ITMs.

* ????

Thurs

* ????

* Check table levelling one last time on all tables. 

Fri

* Close all heavy doors.  (Access connector, ITMX, ITMY, BS, ETMX, ETMY? )

* Drag wipe test masses

* Start at ~10am?

Mon (if not Fri)

* Start pumping

  7339   Tue Sep 4 20:06:04 2012 JenneUpdateLockingMC scan input switched to the 11MHz port of EOM

Since the EOM's signal combiner (splitter backwards) is frequency-independent, Koji and Jamie (in the proper turn off, turn on order) put the 55MHz signal back to the EOM, and put the MC mode scan input to the 11MHz port.  This way we can lock the Michelson tomorrow, and we don't have to keep switching cables around when Riju wants to take some scans.

  7338   Tue Sep 4 20:03:38 2012 JenneUpdateGeneralin-vac progress

Opened ETMY, beam was high and to the right (if you look at the face of ETMY).  Tried walking beam up, since just doing PZT2 caused clipping at the BS before we got to the correct spot on ETMY.  Moved PZT1, then PZT2, to translate beam, but we couldn't get far enough without starting to fall off of PZT2.  Put PZT1 approx. back where it was. Jamie tapped on the top of PR3 (tip tilt just before BS), and then did some compensation with PZT2 to get the beam through the BS target to the center of ETMY.

The beam is very dim at ETMY.  We ended up holding the big IR card with holes such that one of the holes was near the center of the optic, in front of the cage.  Then one person turned off all the room lights so we could see the beam, another person moved PZT2 and PR3, then lights person turned on and off the lights so we could compare beam position with hole position.  A total pain, but it ended up working better than just trying to follow the beam with a card.

We clearly need a better plan for adjusting the tip tilts in pitch, because utilizing their hysteresis is ridiculous.  Koji and Steve are thinking up a set of options, but so far it seems as though all of those options should wait for our next "big" vent.  So for now, we have just done alignment by poking the tip tilt.

Tomorrow, we want to open up the MC doors, open up ETMY, and look to see where the beam is on the optic.  I am concerned that the hysteresis will relax over a long ( >1hour ) time scale, and we'll loose our pointing.  After that, we should touch the table enough to trip the BS, PRM optics, since Koji is concerned that perhaps the tip tilt will move in an earthquake.  Jamie mentioned that he had to poke the tip tilt a pretty reasonable amount to get it to change a noticeable amount at ETMY, so we suspect that an earthquake won't be a problem, but we will check anyway.

After doing all of that, we adjusted IPANG so that the beam gets out of the vacuum envelope.  The beam must squeeze between the wall of the black beam tube and a lens for the oplev, so there is a very, very limited amount of space.  The eventual new tables will be very helpful for alleviating this, but for now we must live with it.  Even though it is pretty squishy right at the edge of the table, with the new layout we should think about giving IPANG a little more space.  Basically, if there is any Yaw motion, the beam going to the QPD will be clipped, and we may get confusing info.  We moved some of the IPANG optics that are on the end transmission table so the beam is centered on the optics while it just makes it through the space between the wall and the lens.  The spot was centered on the IPANG QPD. 

We still need to check on IPPOS, but it is always easier than IPANG.

General notes:

We discovered that the 45deg SOS beam targets are awesome, especially the ones with the irises.  The plain hole ones have very small holes relative to our beam size, so they are much more useful for the MC optics (which they were designed for).  The 45deg targets were made such that the target holes do not line up with the mounting holes.  This is good, since the mounting holes are lower than the center of the optic.  I don't think ericQ and I realized that on Friday, so it's probable that we had installed the target upside-down.  We still need to remake the 0deg targets for the next vent.

Jamie dogged down the new 'bathroom mirror' that lets us see BS and PRM on the same camera view.  He also adjusted some of the masses on the BS table to relevel the table. We need to (at some point) rename the AS_SPARE camera to something like BS_PRM, since we plugged the new camera into the AS_SPARE port on the videoswitch.

tl;dr: Input beam adjusted so we're hitting center of ETMY.  IPANG coming out of vacuum, QPD centered. Need better tip tilt in-situ adjustment capability.

  7337   Tue Sep 4 13:50:26 2012 ericqUpdatePSLPMC Realigned, power adjusted

 I adjusted the PMC alignment this morning, brought the transmission up to 0.83V.

After the lunch meeting, we found the the MC transmission was higher than recently seen. Turned out the HWP had drifted, causing 30mW to be input to the MC. I adjusted it back down to 20mW. 

ELOG V3.1.3-